OpenStudy (anonymous):

I GIVE MEDALS!!!!!!!!!!!!!!!!! Which is an equation in point-slope form of the line that passes through the points (7,5) and (-4,-1)

OpenStudy (studygurl14):

point slope form : \[y - y_{1} = m(x - x_{1})\] \[m = \frac{ y_{2} - y_{1} }{ x_{2} - x_{1} }\] \[m = \frac{ (-1) - (5) }{ (-4) - (7) }\]

OpenStudy (studygurl14):

can you do the rest @sammy34daisy ?

OpenStudy (anonymous):

can u still help not quite sure never learned this

OpenStudy (anonymous):

m= -6/-11

OpenStudy (judygreeneyes):

y = mx + b slope (m) = [5-(-1)]/[7-(-4)] = 6/11 or 0.545 Y = (6/11)x + b plug in one of the points, solve for the intercept -1 = (6/11)(-4) + b b = -1 -24/11 = -35/11 = -3.182

OpenStudy (studygurl14):

I have to go, sorry. Just solve for m, then plus that into the equation as well as y1 and x1

OpenStudy (anonymous):

i dont understand

OpenStudy (studygurl14):

@judygreeneyes , that is slope-intercept form not point-slope

OpenStudy (judygreeneyes):

True -- y – y1 = m(x – x1) is the way you would get the slope, but to solve for m, you still need to rearrange and solve for m = (y-y1)/(x-x1) so to me they end up the same.

OpenStudy (anonymous):

m= -6/-11

OpenStudy (anonymous):



OpenStudy (superdavesuper):

almost there but the term (x-1) is not quite right - check it again plz? :)

OpenStudy (anonymous):

(x+1)

OpenStudy (studygurl14):

y + 5 = 6/11(x + 7)

OpenStudy (superdavesuper):

still no....its for x1 remember...

OpenStudy (anonymous):

(x+4)

OpenStudy (studygurl14):

sorry, "-" not "+" y - 5 = 6/11(x - 7)

OpenStudy (judygreeneyes):

I think you want an x value in the 2nd part y-(-1) = (6/11) (x-(-4))

OpenStudy (superdavesuper):

yup yup u got it...now simplify n u will get the final answer :)

OpenStudy (anonymous):

y+1=6/11(x+4)

OpenStudy (superdavesuper):

questions said slope-point form so u got it :)

OpenStudy (judygreeneyes):

Right, you can use either (x,y) point to find the intercept now that you have the slope. y-5 = (6/11)(x-7) works also

OpenStudy (judygreeneyes):

Good job @sammy34daisy!